# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

### v2

@@ IV-C EquiContact
3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\theta_{1}}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
@@
IV-D Extensions to Pick Tasks
V Experimental Results and Discussions
V-A Benchmark Results
To evaluate the effectiveness of EquiContact and its core design principles, we compare it against three baselines: ACT with world-frame observations and actions, executed with and without GAC, and CompACT [17]. Note that GAC reduces to standard admittance control when the platform is fixed, i.e., not tilted. Table. I summarizes the observation/action representations used in each method and reports the benchmark results across all setups.
Methods Observation Action Cameras Test Scenario Success Rate
ACT w/o GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 2 / 10
ACT w/ GAC [World Pose] [World Pose] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
CompACT [World Pose, FT] [World Pose, Gains] 2 Wrists Cameras Flat Platform (In-Dist.) 20 / 20
Flat Platform (OOD) 0 / 10
EquiContact [GCEV, FT] [Relative Pose, Gains] 2 Wrists Cameras Flat Platform (OOD) 20 / 20
(Place, Ours) Tilted Platform ($30^{\circ}$, OOD) 19 / 20
V-A1 Demonstration of Compliance
As the importance of left-invariant compliant control action has already been verified in [11], we focus on verifying the necessity of compliant control action. We begin by evaluating the role of compliance using the same ACT model architecture, executed with and without the Geometric Admittance Control (GAC). Results for this comparison are shown in the 1st and 2nd rows of Table I. Without GAC, the ACT model shows significantly lower success rates. The failure mode of the ACT w/o GAC involves collision: as the robot approaches the platform, excessive contact forces trigger safety shutdowns, preventing the task from completing. Due to the safety issue, we limited the number of trials without GAC to 10. This result demonstrates that compliant control is nearly a deciding factor between success and failure in contact-rich tasks.
To further analyze the effects of compliant gain control, we compare ACT with GAC and CompACT under in-distribution (In-Dist.) platform settings. Although both methods achieve perfect success rates, their force profiles during insertion differ substantially. As shown in Fig. 4, CompACT, which outputs task-adaptive admittance gains based on force-torque feedback, consistently produces lower interaction forces in all directions. The effectiveness of the CompACT compared to the baseline ACT was already presented in [17].
V-A2 Demonstration of Equivariance
Although the CompACT succeeds in insertion tasks in trained scenarios without excessive force exertion, it fails to generalize to spatially unseen configurations. This is expected, as its observation and action representations are defined in the global spatial frame, which neither guarantees nor encourages equivariance. The result of applying CompACT to the translationally unseen cases is shown in the Table I 33rd row - Flat Platform (OOD). Note that the flat platform is only utilized, meaning it is only randomized translationally. We tested for $10$ cases and did not conduct more tests because it resulted in a $0\%$ success rate.
TABLE II: Success Rates of the proposed EquiContact for a full pipeline of pick and place.
Test Scenario Success Rate Failure Cases
Flat Platform (OOD) 20 / 20 N/A
Tilted Platform ($30^{\circ}$, OOD) 18 / 20 1 Pick, 1 Place
The result of EquiContact for the full pick-and-place task is summarized in Table II. The EquiContact also demonstrates a near-perfect success rate in the full pick-and-place pipeline for peg-in-hole tasks. However, as the whole task is formulated in sequential stages, the error in the previous stage tends to propagate to the following stages, leading to slightly increased failure cases.
V-B Limitations and Discussions
TABLE III: Success Rates of the G-CompACT (insertion) trained with the base dataset and augmented dataset for the flat platform under the presence of visual distractor scenario and tilted platform with large angle cases. The evaluation is conducted with the ground-truth reference frames.
Test Scenario Base Dataset Aug. Dataset
W/ Visual Distractor 4 / 10 9 / 10
Tilted Platform ($45^{\circ}$) 5 / 10 9 / 10
V-B1 Limitation on Left-invariant Visual Features
Although EquiContact is theoretically vision-to-force equivariant, this guarantee holds only under Assumptions 1 and 2. In particular, Assumption 2, which requires left-invariant visual features, is challenging to ensure in practice, as our approach does not explicitly encode this property through loss functions or architectural inductive bias. Consequently, G-CompACT shows degraded performance in settings with visual distractors or severe tilting ($45^{\circ}$), as shown in Table III (“Base Dataset”). The degradation arises because larger tilt angles expose unseen background regions in the wrist camera view (see Fig. 3), resulting in out-of-distribution observations. This can be mitigated by augmenting the training data with visual distractors and demonstrations collected from a tilted angle. To demonstrate this, we added $20$ demos with distractors on a flat platform and 20 demos from a $30^{\circ}$ tilted setup. As shown in the “Aug. Dataset” column of Table III, the success rates have recovered to original levels.
Enforcing left-invariance on RGB images via architectural design remains an open challenge. Recent works have explored inducing 3D equivariance from 2D images [32, 33]. Another promising direction involves using pretrained foundation models, such as CLIP and DINO, which may exhibit implicit equivariance due to large-scale and diverse training. Language-guided features (e.g., using the prompt “peg”) can help extract relevant features across views. Alternatively, real-time segmentation models can be used to focus only on the object of interest. We will investigate enforcing left-invariance of images in future work.
V-B2 Hardware and Algorithmic Limitation
A key challenge in implementing equivariant policies for robotic manipulation is handling singularities [10]. When the robot is near singular, pose tracking accuracy degrades because controllers sacrifice tracking to avoid singularity, leading to poorly executed policy commands. This results in a distributional shift, which is a critical limitation for imitation learning. From our experimental demonstration perspective, a small workspace of the Diff-EDF (Fig. 2) was a drawback due to its limited receptive field. A combination of the singularity, joint limits, and small workspace resulted in limited demonstration of our approach, not covering a complete $SE(3)$ transformation.
V-C Validating Feasibility to Other Contact-rich Tasks
To further validate the EquiContact framework, we test it on two additional contact-rich tasks: screwing and surface wiping (see Fig. 5). In the screwing task, the robot aligns and screws a peg; in the wiping task, it erases a line from a whiteboard. For simplicity, we use a white background to eliminate visual distractions and assume that reference frames are known. The reference frame for screwing is the end-effector pose at full insertion; for wiping, it is the center of the board. As in the PiH setup, demonstrations are collected on a fixed platform and evaluated on out-of-distribution configurations, including tilted platforms. We collected 50 demonstrations per task. Results in Table IV show consistent success rates across all conditions, confirming EquiContact’s spatial generalizability.
Figure 5: (Left) Screwing task, (Right) Surface wiping (erasing) task. The same platform structure of the PiH is used.
TABLE IV: Success Rates of the G-CompACT for screwing and surface wiping tasks. The evaluation is conducted with the ground-truth reference frames. Both tasks are trained with $50$ demos.
Test Scenario Screwing Surface Wiping
Flat Platform (OOD) 10 / 10 10 / 10
Tilted Platform ($30^{\circ}$, OOD) 9 / 10 9 / 10
VI Conclusion
In this work, we introduced EquiContact, a vision-to-force equivariant policy for spatially generalizable contact-rich tasks. By integrating a global reference frame estimator (Diff-EDF) with a fully localized visuomotor servoing policy module (G-CompACT), we demonstrate how compliance, localized policy, and induced equivariance can be unified to enable the peg-in-hole (PiH) task, a representative contact-rich precision task, under spatial perturbations. We proved the $SE(3)$ equivariance of the policy under assumptions on point cloud and image observations, validated its effectiveness through real-world experiments on PiH benchmarks and its feasibility towards screwing and surface wiping tasks. Compared to benchmark methods, our approach generalizes to unseen platform positions and orientations while maintaining low contact force and near-perfect success rates. Through extensive benchmark studies, we highlighted the effectiveness of the three principles – compliance, localized policy, and induced equivariance – for achieving spatial generalizability in contact-rich manipulation. We conclude that these principles offer a simple yet powerful design guideline for developing spatially generalizable and interpretable robotic policies complementing recent trends in end-to-end visuomotor learning and enabling a structured divide-and-conquer approach.
Supplementary Material
Implementation Details
In this section, we provide the details of the hardware code implementation. Since every hardware system has a different implementation, our goal is to provide a comprehensive overview and a guide to our code distribution. We will use the exact names of the nodes to enable readers to understand our codebase easily. The hardware experimental implementation code is provided in https://github.com/Joohwan-Seo/EquiContact-Experiment, and the simulation implementation code in https://github.com/Joohwan-Seo/EquiContact-Simulation.
S-A Implementation Pipeline
The pipeline visualized in Fig. S1 is an interconnected system of camera, robot, and controller nodes with two policy services, all implemented in ROS2 Humble. The two policies are implemented as services because they are activated infrequently and latencies in activating the policies initially do not affect the behavior in practice. The execution of the full pipeline is divided into two primary stages: querying Diffusion-EDF for initial pick and insertion end-effector poses, and then executing the appropriate ACT sub-policies for these two steps after reaching a neighborhood of the predicted poses. In the first stage, a request to the Diffusion-EDF Service is made through the FullPipelineClient, which prompts the service to pull the individual point clouds published by the Orbbec Node. The service subsequently merges the point cloud in the spatial frame, preprocesses it, and passes it to the initialized Diff-EDF model to produce both pick and insertion poses. The Diff-EDF model, as covered in the following section, is formulated as a “double-pick” model that assumes that a fixed grip will be executed, and a cached empty gripper point cloud is passed along with the scene point cloud to get both predictions.
In the second stage, the FullPipelineClient requests the G-CompACT Service node to complete the task with these two poses as noisy references. The service initializes both pick and place policies and executes the tasks serially, raising flags to signal completion of each step. For raw camera images and proprioceptive information, the service subscribes to nodes for each camera (from Wrist Camera Nodes) and the Robot Node respectively; note that the controller and proprioceptive helper are encapsulated in one Robot Node that manages both publishing proprioceptive information and executing policy predictions. During the pick portion, the pick ACT model only publishes end-effector poses, and fixed high admittance gains are utilized because the demonstrations encourage the model to make limited contact with the peg. The pick model has a fixed time to complete the task before a grasp is attempted, and the robot moves to the next step. Similarly, the place ACT policy is executed given a fixed-time limit, but for the place model, the force-torque data is also passed as an input, and admittance gains are outputted along with the end-effector poses. A fixed-time limit for both phases is used because it is easily implemented and works more consistently than other, more complicated completion heuristics.
Figure S1: Whole pipeline implemented (via client named as FullPipelineClient) with ROS2 is presented.
S-B Diffusion-EDF
S-B1 Implementation Heuristics
We also used the post-processing heuristics to filter the output pose of the Diff-EDF. The Diff-EDF outputs $20$ candidate target poses for picking and $10$ candidates for the place, which are ranked by the energy level. Although in theory, the lower energy poses should result in a better pose, we found out that this does not hold in practice. Instead, we figured out that the Diff-EDF have low variations on the position of the “tip”. To elaborate, suppose a picking policy returns the desired poses $g_{pick}^{i}=(p_{pick}^{i},R_{pick}^{i})$, which are the end-effector poses, not the tip, or Tool-Center-Point (TCP) poses. Then, the position of the tip estimated by Diff-EDF can be calculated via
$p_{tip,pick}^{i}=p_{pick}^{i}+R_{pick}^{i}p_{et,pick}$
where $p_{et,pick}\in\mathbb{R}^{3}$ denotes the coordinate of the gripper’s tip with respect to the end-effector body frame. Using this property and the task setup that the peg is always upright, which gives $R_{peg}$, we recalculate the Diff-EDF of the picking as
$p_{pick}=\bar{p}_{tip,pick}-R_{peg}^{T}p_{et,pick}$
where $\bar{p}_{tip,pick}$ is the mean of $p_{tip,pick}^{i}$ for $i=1,\cdots,N$. On the other hand, for placement, we did not make any assumptions about the orientation of the hole-platform structure. The heuristic we used for the place is as follows:
$p_{place}=\bar{p}_{tip,place}-\bar{R}_{place}^{T}p_{et,place},$
where $\bar{R}_{place}$ is the averaged rotation of $R_{place}^{i}$, and $\bar{p}_{tip,place}$ is the averaged position of the tip calculated similarly from the picking case.
S-C Tips on applying Geometric Admittance Control
Here, we present a few tips on the application of geometric admittance control. The GAC is implemented as a part of a ROS2 node, running with $200\mathrm{Hz}$.
S-C1 ROS2 Implementation
The Robot Node acts as the central controller, receiving and executing desired pose commands using a Geometric Admittance Control (GAC). Its main execution loop operates at $200\mathrm{Hz}$, ensuring smooth and compliant motion in response to external interactions. The node subscribes to the “action” topic, which contains the pose command and admittance gains command. Additionally, the controller subscribes to the “rebias” topic in order to preserve the accuracy of force readings. When a message on this topic is received, a callback raises a flag, triggering a re-biasing procedure for the FT sensor during the next control cycle. Within its main loop, the controller sends the intended pose and processed FT sensor data to the GAC function, which generates the compliant commands for the robot’s low-level motion controller.
S-C2 Handling Force-Torque Sensor Data
The first consideration is how to handle the Force-Torque (FT) sensor attached to the end-effector. It is always favorable to subtract the bias of the FT sensor after finishing each task, such as picking up the peg. We collect the FT sensor value for $2$ seconds and take a mean value to use as a bias. The other is to filter the FT sensor. Since the FT sensor is noisy, it is strongly recommended to implement a low-pass filter. However, having a low cut-off frequency can lead to serious phase lag, resulting in so-called “bouncing” behavior. We designed a simple 1st-order low-pass filter and implemented it using a state-space representation in discrete time.
S-C3 Implementation Tips
As mentioned in Section. IV, the detailed implementation method depends on the low-level servo controller type. Typical controllers provided by robot manufacturers are joint position control, joint velocity control, end-effector position control, and end-effector velocity control. If a torque controller is readily available, then directly implementing a geometric impedance controller (GIC) is recommended, but joint friction compensation needs to be handled appropriately. Given the desired dynamics (9), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (16)
where $T_{s}$ is a sampling time ($5\mathrm{ms}$ in our case for GAC) and $\hat{(\cdot)}$ denotes a hat-map.
The end-effector velocity control can be straightforwardly applied from (S-C3), by directly using $V_{d}^{b}(k)$ as a desired end-effector velocity command. One caveat here is the $V^{b}$, which is body-frame end-effector velocity. The common velocity provided by the manufacturer is defined on the Cartesian (Spatial) frame. If end-effector velocity is only offered in the Cartesian frame $V\in\mathbb{R}^{6}$, then
$V^{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}V,$
where $R$ is a rotation matrix from the current configuration. However, sometimes, it is provided in the Cartesian frame for translational velocity and the body frame for rotational velocity. The users should verify the frames of the velocities by running the robot in one direction and logging its values. For the joint space-based approach, one can use
$\begin{split}\dot{q}_{d}(k)\!&=\!\dot{q}(k)\!+\!T_{s}\cdot J_{b}(k)^{-1}M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),\\ q_{d}(k)\!&=\!q(k)+T_{s}\cdot\dot{q}_{d}(k),\end{split}$
where $J_{b}$ denotes a body-frame Jacobian matrix. $\dot{q}_{d}(k)$ is provided to joint-velocity controller and $q_{d}(k)$ is provided to the joint-position controller. The Jacobian matrix also needs to be identified concerning the end-effector velocity, i.e., if the end-effector velocity is given in the Cartesian frame, then the Jacobian matrix $J$ is also provided in the Cartesian frame. In this case, the body-frame Jacobian matrix can be obtained via
$J_{b}=\begin{bmatrix}R^{T}&0_{3\times 3}\\ 0_{3\times 3}&R^{T}\end{bmatrix}J$
When the Jacobian matrix is not provided, an alternative is to use the robotics toolbox from Python or Matlab. First, use the DH parameters provided by the manufacturer (or URDF), then build the Jacobian matrix function. The function can be cast into a lambda function in Python. For Python, the Jacobian matrix function can also be wrapped into executable .so file using Matlab code builder and executed in the loop for fast computation (well below $1\mathrm{ms}$).
S-D Teleoperation Details
The architecture of our teleoperation system is explained to make it easier to reproduce and adapt our work. The system is constructed modularly, with the human-operator interface decoupled from the robot controller. This allows greater adaptation to different robotic platforms and control algorithms. The teleoperation function is implemented by two ROS nodes: the $\mathtt{RobotNode}$ (named as $\mathtt{robot\_gac}$ in the code implementation), which processes proprioceptive information and executes a Geometric Admittance Controller to carry out operations, and the $\mathtt{Teleoperation}$ (named as $\mathtt{teleop\_ros}$ in the code implementation) node, which handles operator input.
The teleoperation node acts as the principal interface between the human operator and the robot controller. A 6-DOF SpaceMouse is used to record input, which is then converted into the necessary end-effector pose. The main teleop function of this node operates at a frequency of $200\mathrm{Hz}$ inside a timer loop. It adjusts the desired end-effector pose incrementally during each cycle by reading velocity commands from the SpaceMouse.
The SpaceMouse provides velocity commands $V_{sm}=[v_{sm}^{T},\omega_{sm}^{T}]^{T}\in\mathbb{R}^{6}$, where $v_{sm}\in\mathbb{R}^{3}$ is translational velocity and $\omega_{sm}\in\mathbb{R}^{3}$ is rotational velocity. The change in desired position, $\Delta p$, is computed by scaling $v_{sm}$ with a position offset:
$\Delta p=v_{sm}\cdot\mathtt{pos\_offset}$
The target position is then updated as:
$p_{d}(k+1)=p_{d}(k)+\Delta p$
The desired end-effector orientation, $R_{d}(k+1)$, is updated by composing the previous rotation with the incremental rotation matrix, $\Delta R$. The small, incremental angles, $\Delta\theta$, are computed by scaling $w_{sm}$ with rotation offset:
$\Delta\theta=w_{sm}\cdot\mathtt{rot\_offset}$
The $\Delta\theta$ is then converted to $\Delta R$ using Euler angle as $\Delta R=\mathtt{euler2rotmat(\Delta\theta)}$. The target orientation is then updated as:
$R_{d}(k+1)=R_{d}(k)\cdot\Delta R$
This calculated pose is then published as a command for the robot controller at a frequency of $200\mathrm{Hz}$. To create a dataset of the teleoperation commands, a separate timer logs the same message at $30\mathrm{Hz}$. This implements a teleoperation that operates on the robot base frame for translational movement and the end-effector frame for rotational movement.
This teleoperation node was created with multiple operation modes to improve usability. For end-effector alignment with tilted platforms, the operator can switch the SpaceMouse’s control frame between the end-effector frame and the robot’s base frame. Additionally, a dedicated button allows users to modify the sensitivity offsets for position and rotation, enabling users to alternate between a translation-focused and an orientation-focused mode. The operator can globally scale motion sensitivity by cycling through low, medium, and high speed levels. The system also includes specific features, such as an “insertion mode” that operates in the end-effector frame for the translational movement to help with high-precision alignment tasks, as well as requests to set the robot’s end effector at “home” or a randomized pose.
Furthermore, the operator can dynamically adjust the parameters of the Geometric Admittance Controller via dedicated keyboard inputs. There are four gain profiles: Free mode applies high stiffness gains $(1000)$ to all axes for precise motion in open space; Contact mode uses low z-axis gain $(300)$ with high gain elsewhere $(1500)$ to exert less force on surfaces; Insertion mode reverses this, with high z-axis gain $(1500)$ for precision while allowing compliance in other directions $(300)$; and Compliant mode sets all gains low $(300)$ to make the end-effector highly responsive to external forces. These gain profiles are published as action messages, together with poses and gripper states.
S-E G-CompACT Training Details
During the training of the G-CompACT, we added the noise to the known fixed pose of the hole and peg. As mentioned earlier, this is because the output of the Diff-EDFs $g_{{}_{EDF}}$ is biased. By adding the noise to the goal pose signal, we encourage the model to focus on the image of the wrist camera instead of the incorrect GCEV $e_{{}_{G}}$ signal. For our training, we added noise to every step of the demonstration dataset drawn from a uniform distribution. The noisy target poses $\tilde{g}_{ref}=(\tilde{p}_{ref},\tilde{R}_{ref})$are calculated as
$\begin{split}\tilde{p}_{ref}&=p_{ref}+n_{p},\\ \tilde{R}_{ref}&=R_{ref}\cdot\mathtt{euler2rotmat}(n_{r}),\end{split}$
where $n_{p}\sim\mathcal{U}(-0.02,0.02)\in\mathbb{R}^{3}$ and $n_{r}\sim\mathcal{U}(-8^{\circ},8^{\circ})\in\mathbb{R}^{3}$, which denotes $2\mathrm{cm}$ of translational and $8^{\circ}$ (degree) of rotational error. The impact of adding noise to the reference frame is further shown in Appendix. -C.
To encode the element on $SE(3)$ for learning, we used position & rotation vector representation for G-CompACT. However, in benchmark models of CompACT and ACTs, we utilized position & rot6d representation. This is because the rotation vector representation has a sign flip issue when the rotation angle is near $180^{\circ}$, but the rot6d representation does not have that problem. For G-CompACT, as the actions are represented with respect to the current end-effector frame, the transformation angles are relatively small and do not have a sign flip problem.

### v3

@@ VI Conclusion
3:Run Diff-EDF for reference frame $g_{{}_{EDF}}\!=\!f_{\varphi}(\mathcal{O}^{scene},\mathcal{O}^{grasp})$
@@
Figure A2: Whole pipeline implemented with ROS2 is presented.
As mentioned earlier, the full EquiContact pipeline is implemented with ROS2 framework. The whole implementation flow is presented in Fig. A2, and also summarized as in Algorithm 1.
AIII Additional Experimental Results
AIII-A Errors of Diff-EDF
The RMSE error of the Diff-EDF on the training dataset is presented in Table A2. The RMSE of the rotational error is naively calculated from the Euler angles of the error rotation matrix.
TABLE A2: RMSE error values of Diff-EDFs on the training dataset. The dimensions of translational errors in $x,y,z$ directions, given by $e_{T,x},e_{T,y},e_{T,z}$, are $\mathrm{mm}$ and the rotational errors in $x,y,z$ directions, given by $e_{R,x},e_{R,y},e_{R,z}$, are $\mathrm{deg}$.
$e_{T,x}$ $e_{T,y}$ $e_{T,z}$ $e_{R,x}$ $e_{R,y}$ $e_{R,z}$
pick 7.173 6.933 6.199 7.650 15.90 15.67
place 13.75 8.241 5.999 3.806 5.560 5.660
As noticed from the table, the translational error is significantly larger than the desired accuracy of precision of the PiH task $\sim 1\mathrm{mm}$. In addition, the rotational error of the picking task is significantly higher than that of the placing task. Therefore, we use the “upright peg” assumption for the full pipeline implementation.
AIII-B Vision Encoder Design Study
Here, we conduct a controlled comparison of vision encoder variants. To verify the design choices to meet the conditions of Assumption 1, we have trained $4$ models with the same training dataset for PiH tasks, which are listed below:
  • Baseline ACT architecture that uses ResNet 18 and without language feature (RN18)
  • ACT with pretrained CLIP-RN50 but is frozen (CLIP-RN50-frozen)
  • ACT with CLIP-RN50, but $10\%$ of learning rate for vision backbone (CLIP-RN50-SB, SB stands for slow backbone training)
  • ACT with CLIP-RN50, same learning rate for policy and vision backbone (proposed)
We have tested our models in the in-distribution condition and with a $45^{\circ}$ transformation in the $y$ axis, i.e., the third case for extreme task transformations (Fig. 3). The results are summarized in Table A3.
TABLE A3: Results of vision encoder design study. OOD case here is a $45^{\circ}$ transformation in the $y$ axis.
Backbone Learning Rate Learning Rate Success Rate
($\eta_{policy}$) ($\eta_{vision}$) In-dist OOD
RN18 1e051e-05 1e051e-05 10 / 10 6 / 10
CLIP-RN50-frozen 1e051e-05 0 3 / 10 3 / 10
CLIP-RN50-SB 1e051e-05 1e061e-06 10 / 10 0 / 10
CLIP-RN50 (proposed) 1e051e-05 1e051e-05 10 / 10 10 / 10
We first observe that the vision encoder without language guidance degrades under the OOD rotation (6/10), although background randomization during data collection partially mitigates background overfitting. In contrast, using a frozen CLIP-RN50 encoder yields low success even in-distribution (3/10), suggesting a significant domain mismatch between internet-scale pretraining and the short-range wrist-camera viewpoint in contact-rich manipulation. Interestingly, fine-tuning the CLIP-RN50 encoder with a very small learning rate achieves high in-distribution performance (10/10) but fails completely under the OOD rotation (0/10). We speculate that the visual representation is not sufficiently adapted: the encoder adjusts only locally to the training task configuration, without acquiring robustness to large geometric shifts, making the downstream policy brittle when viewpoint changes substantially. Finally, jointly fine-tuning the CLIP-RN50 encoder together with the policy (proposed) recovers both in-distribution and OOD performance (10/10), indicating that stronger encoder adaptation is critical for wrist-camera generalization under task transformations.